If a jury consists of 10 men and 15 women, what is the probability of selecting a 6-person jury consisting of all females?
Found 2 solutions by oscargut, nerdybill:
Answer by oscargut(2103)

Solution:
P(6 females) = C(15,6)/C(25,6)= 5005/177100 = 0.0283
If you have any doubt or if you have more problems my e-mail is:
mthman@gmail.com

Answer by nerdybill(7384)

If a jury consists of 10 men and 15 women, what is the probability of selecting a 6-person jury consisting of all females?
.
"number of ways to select 6-females from 15"/"number of ways to select any 6 from the total group":
.
total number of people to pick from: 10+15=25
.
15C6/25C6
5005/177100
.028
2.8%